ON COVERING TRANSLATIONS AND HOMEOTOPY GROUPS OF CONTRACTIBLE OPEN n-MANIFOLDS

ROBERT MYERS

Abstract. This paper gives a new proof of a result of Geoghegan and Mihalik which states that whenever a contractible open n-manifold W which is not homeomorphic to R^{n} is a covering space of an n-manifold M and either n ≥ 4 or n = 3 and W is irreducible, then the group of covering translations injects into the homeotopy group of W.